# NoiseBench: Benchmarking the Impact of Real Label Noise on Named Entity Recognition

### v1

@@ Abstract
Available training data for named entity recognition (NER) often contains a significant percentage of incorrect labels for entity types and entity boundaries. Such label noise poses challenges for supervised learning and may significantly deteriorate model quality. To address this, prior work proposed various noise-robust learning approaches capable of learning from data with partially incorrect labels. These approaches are typically evaluated using simulated noise where the labels in a clean dataset are automatically corrupted. However, as we show in this paper, this leads to unrealistic noise that is far easier to handle than real noise caused by human error or semi-automatic annotation. To enable the study of the impact of various types of real noise, we introduce NoiseBench, an NER benchmark consisting of clean training data corrupted with 6 types of real noise, including expert errors, crowdsourcing errors, automatic annotation errors and LLM errors. We present an analysis that shows that real noise is significantly more challenging than simulated noise, and show that current state-of-the-art models for noise-robust learning fall far short of their theoretically achievable upper bound. We release NoiseBench to the research community [^1] https://github.com/elenamer/NoiseBench.
@@
Elena Merdjanovska1,2, Ansar Aynetdinov1, Alan Akbik1,2 1Humboldt University of Berlin 2Science of Intelligence {elena.merdjanovska, aynetdia, alan.akbik}@hu-berlin.de
@@ 1 Introduction
Named entity recognition (NER) is the task of detecting and classifying named entities in text, such as the names of organizations or locations. All current state-of-the-art approaches for NER require supervision in the form of labeled training data, i.e. sentences in which named entities are marked and assigned their correct type. However, prior work found that available datasets for NER and other supervised tasks are affected by label noise, meaning that a certain percentage of entity labels are incorrect. For instance, the common NER dataset CoNLL-03 Tjong Kim Sang and De Meulder (2003) was estimated in various studies to have noise shares of between 5 and 7% Wang et al. (2019); Reiss et al. (2020); Rücker and Akbik (2023). Other NER datasets have also been found to contain a share of incorrect labels, with Ontonotes4 estimated around 8% and WNUT-17 around 18% Wang et al. (2019); Huang et al. (2021).
@@
  • We comparatively evaluate current state-of-the-art approaches for noise-robust learning on NoiseBench, and experimentally establish theoretical upper bounds.
@@ 2 NoiseBench
Our benchmark is derived from a subset of the classic CoNLL-03 dataset for NER in English, annotated with entities belonging to four classes. We chose this dataset since it has been extensively studied in the field, allowing us to integrate various prior works.
NoiseBench consists of the following parts: (1) A noise-free test split to evaluate trained models. (2) Seven variants of the training split, where six are annotated with different types of noise and one is without noise. The quality of the six noisy variants, in reference to the noise-free dataset, is presented in Table 2.
@@
Table 1: Overview of the noisy training splits in NoiseBench. The table shows the noise level, the micro-averaged token-level F1 score (F1 token ), micro-averaged entity-level F1 (F1 entity ), the number of entities (Total), number of correct entities (Correct) and share of each error type: missing mentions (Missing (FN)), labelled non-entity mentions (Non-entity (FN)) and wrong type (Type), as well as the number of partial matches (Partial). All metrics are in comparison to the Clean split.
@@ 2.1.2 Expert Errors
The largest portion of machine learning datasets relies on manual annotation by domain experts to provide high-quality labels. However, errors have been found to occur even in expert annotation, affecting even well-known benchmarks, though usually with relatively low noise shares of under 10% Northcutt et al. (2021b); Song et al. (2022). To represent such noise, our benchmark includes a variant of the train split called Expert, which contains the original CoNLL-03 annotations. As Table 1 shows, this split has a noise share of 5.5% and is thus the split with lowest noise.
@@ 2.1.3 Crowdsourcing Errors
In addition to expert annotations, labeled datasets can be obtained via crowdsourcing. These annotations are heavily prone to human annotation errors, which can happen for different reasons (Frenay and Verleysen, 2014). In order to create noisy variants of the train set in our benchmark representing real-world human errors, we utilize the crowdsourced labels by Rodrigues et al. (2014). This study involves 47 crowd workers labelling a subset of the CoNLL-03 dataset, of around 400 news articles. They released their dataset and all annotations produced by each crowd worker. We selected only the sentences where the tokenization matched the Clean variant, resulting in 5,885 sentences.
@@ 2.1.4 Distant Supervision
One approach for labeling data without human participation is distant supervision Mintz et al. (2009), where entity mentions in target datasets are matched to entity types in knowledge bases (KBs). The matching can be achieved by simple string matching, use of regular expressions or heuristics.
We include a Distant noisy training variant in our benchmark, adapted from the annotations by Liang et al. (2020) [^2] Available under Apache 2.0 license that use the Wikidata corpus and gazetteers collected from multiple online sources as external knowledge bases. After initial POS tagging, the unlabeled sentences were matched with the knowledge bases.
This process results in incomplete annotations due to limited coverage over entity types of KBs. This explains the rather high number of missing entities and the overall noise level (31.3%) of the Distant training variant, as shown in Table 1.
@@ 2.1.5 Weak Supervision
Another approach aimed at reducing manual annotation efforts is weak supervision. Here, labels are obtained using a number of ’weak’ supervision sources, such as heuristics or expression-based rules. Each weak source is typically specialised to detect only a subset of the correct labels.
We use the labels from the approach by Lison et al. (2020)2 to create our Weak label set. This covers 16 weak labeling sources, those used in Zhang et al. (2021b), including heuristics, gazetteers and predictions of NER models trained on other corpora. An example heuristic is detecting PER (person) entities using a pre-defined list of first names.
@@ 2.1.6 LLM Teacher Models
Our benchmark includes a noisy variant of the train split annotated by an LLM. This follows recent efforts that use LLMs for dataset generation Wang et al. (2023). Here, the main idea is to pass a natural language description of the annotation task and target classes to an LLM, and provide sentences that the LLM should label. Current works find that LLMs are able to generate high quality labels for some tasks (e.g. sentiment classification) while for others (e.g. NER and question type categorization) the resulting labels are very noisy.
@@ 2.2 Statistics
Furthermore, the table shows the total number of entities, the number of correct entities, as well as the share of different error types. The errors are categorized into 4 main categories: missing mentions, non-entity mentions (false positives), incorrect entity type (where the boundary is correct, but type incorrect) and partial matches. Partial matches are special cases where the type is correct, however the mention boundary is only partially correct. Refer to Figure 1 for examples.
@@ 3 Comparing Real and Simulated Noise
3.1 Noise Simulation Methods
3.2 Experimental Setup
Validation splits. We evaluate the setting in which all available data to train a model is noisy, including the validation set.
To obtain noisy validation sets for each of our 7 dataset variants, we split the noisy datasets into training and validation sets. All sentences from 66 news documents from 1996-08-24 comprise the validation set, which is around 17% of all sentences, and are left out from model training and used for hyperparameter tuning.
Baseline. For NER, as a baseline approach, we fine-tune an xlm-roberta-large transformer using the FLERT approach Schweter and Akbik (2021). It improves upon the regular fine-tuning setup by considering document-level features of a sentence to be tagged. We use a learning rate of 5e-6 and a batch size of 32, for a fixed number of 10 epochs. These parameters were obtained according to the performance on a noisy validation set, keeping in mind that larger batch sizes are more robust to noise Rolnick et al. (2017). As an evaluation metric we use entity-level micro F1-score.
@@ 3.2 Experimental Setup
Figure 2: Comparison of model performance during extended training. The top row shows models fine-tuned on label sets with real noise, while the bottom row models fine-tuned on a corresponding simulated (class-dependent) noisy label set. The graphs show the F1 score on the noisy training labels and on the clean training labels, for 3 noise types. The plots are averages of 3 runs. The remaining 3 noise types are shown in Appendix C.
3.3 Experiment 1: Impact of Label Noise on Test Performance
3.3.1 Results and Discussion
The results for uniform noise are shown in Appendix A. We initially established that uniform noise is significantly less challenging for the model, so in the results from Experiments 1 and 2 we chose to focus our analysis solely on oracle class-dependent noise.
The main results from Experiment 1 for oracle class-dependent noise are shown in Table 2. Additional analysis of the results can be found in Appendix B. Following are our main observations.
Label noise degrades performance. When we compare the test F1 scores of the real noisy variants with the average score of 93.99 achieved when training on the Clean variant, we can see that model performance is affected by each noise type. In all cases, the noisy training variants yield lower test scores, even when training with the low-noise Expert variant. As the noise levels increase, the impact on model performance becomes more pronounced. This shows that the baseline model lacks robustness to any of the real noise types. When we compare the test F1 scores of the simulated noisy variants, we can see that noise of 5.9 in the training set does not hurt model performance and results in a score comparable to training on the Clean variant. However, as simulated noise levels increase, they do degrade the prediction scores on the test set.
@@ 3.4.1 Results and Discussion
In Table 2 we show training curves from training with real and simulated variants of NoiseBench  for 3 noise types: Expert, Crowd++ and Distant. We plot two scores: the F1-score on the respective noisy variant of the training set, and the F1 score on the Clean variant of the training set. In all training curves, we can observe the memorization effect, with each model perfectly fitting the noisy data by the end of training and reaching an F1 score close to 1.
Delayed memorization of simulated noise. However, we note that with simulated noise (see Table 4, 4, 4) this happens much later in the training process than with real noise. In addition, the training curves of simulated noise show a stage during the early epochs where the score on the clean labels is consistently higher than the score on the noisy labels. This confirms previous findings that the model is able to learn general patterns first, before starting to memorize the noise.
Immediate memorization of real noise. With real noise this does not happen and the model starts fitting the noisy labels from the beginning of training (see Table 4, 4, 4). As a result, the score on the clean labels is consistently lower than the score on the noisy labels, during the entire training run.
Our experiments find that real noise does not display distinct generalization/memorization phases during training, and rather immediately begins with memorization. This makes intuitive sense, as real noise has underlying patterns that may be extracted during learning. This lends further evidence to the increased challenges and the need to evaluate noise-robust learning with real noise.
@@ 4.1 Compared Approaches
We surveyed current state-of-the-art methods for noise-robust NER and found that many approaches rely on the same underlying ideas for handling label noise. In the following, we group approaches by the underlying idea, select a state-of-the-art representative for each group and, if possible, derive an upper bound method for each noise-robust method. For more details about the implementation of compared approaches refer to Appendix D.
@@ 4.1.1 Learning from a Clean Subset
The first family of approaches relies on utilizing the subset of each noisy dataset in which all labels are correct. They either filter out all likely incorrect annotations and learn only from a clean subset Northcutt et al. (2021a). Or they derive confidence weights for each annotation so that annotations judged to be of higher quality feature more during training, as in the CrossWeigh and COSINE approaches Wang et al. (2019); Yu et al. (2021). As representative of this class of approaches, we chose confident learning Northcutt et al. (2021a).
Upper bound: Oracle subset. To obtain an upper bound for this family of approaches, we use an oracle to select the subset of clean sentences from each of the noisy training splits in NoiseBench. We then use a standard fine-tuning approach only on this subset. This setting illustrates the best-case scenario for training on a clean subset only.
@@ 4.1.2 Delaying Memorization
Another family of noise-robust learning approaches seeks to leverage the two phases of learning (generalization and memorization) we discussed in Section 3.4. They seek to either draw our the generalization phase or cease training before memorization begins. While our experiments indicate that these two phases do not exist for real noise, we nevertheless include this family of approaches in our evaluation since they are widely used and to gain additional insights into their behavior on real noise. As representative of this class of approaches, we chose co-regularization (Zhou and Chen, 2021).
Upper bound: Oracle stopping. To obtain an upper bound for this family of approaches, we use a simple stopping criterion based on the score on the clean test set at the end of each epoch. We use the epoch of best generalization to report the final score. This simulates an ideal stopping (albeit at the granularity of full epochs).
@@ 4.1.3 Combined Approaches
While the approach discussed so far each build on the individual ideas of identifying a clean subset or delaying memorization, many current approaches in fact combine multiple of such ideas in multi-stage pipelines Liang et al. (2020); Yu et al. (2021); Wang et al. (2022). As representative of such approaches, we evaluate BOND Liang et al. (2020), which combines pseudo-labeling in a student-teacher setup and sample selection.
No upper bound for pseudo-labeling. We cannot derive a separate upper bound for pseudo-labeling, as the best case scenario here would mean that all noisy labels are replaced by correct labels. This means that the upper bound for pseudo-labeling is the same as training on fully clean data.

### v2

@@ Abstract
Available training data for named entity recognition (NER) often contains a significant percentage of incorrect labels for entity types and entity boundaries. Such label noise poses challenges for supervised learning and may significantly deteriorate model quality. To address this, prior work proposed various noise-robust learning approaches capable of learning from data with partially incorrect labels. These approaches are typically evaluated using simulated noise where the labels in a clean dataset are automatically corrupted. However, as we show in this paper, this leads to unrealistic noise that is far easier to handle than real noise caused by human error or semi-automatic annotation. To enable the study of the impact of various types of real noise, we introduce NoiseBench, an NER benchmark consisting of clean training data corrupted with 6 types of real noise, including expert errors, crowdsourcing errors, automatic annotation errors and LLM errors. We present an analysis that shows that real noise is significantly more challenging than simulated noise, and show that current state-of-the-art models for noise-robust learning fall far short of their achievable upper bound. We release NoiseBench for both English and German to the research community [^1] https://github.com/elenamer/NoiseBench.
@@
Elena Merdjanovska1,2, Ansar Aynetdinov1, Alan Akbik1,2 1Humboldt-Universität zu Berlin 2Science of Intelligence {elena.merdjanovska, aynetdia, alan.akbik}@hu-berlin.de
@@ 1 Introduction
Named entity recognition (NER) is the task of detecting and classifying named entities in text, such as the names of organizations or locations. Current state-of-the-art approaches for NER still require supervision in the form of labeled training data Zaratiana et al. (2023), i.e. sentences in which named entities are marked and assigned their correct type. However, prior work found that available datasets for NER and other supervised tasks are affected by label noise, meaning that a certain percentage of entity labels are incorrect. For instance, the common NER dataset CoNLL-03 Tjong Kim Sang and De Meulder (2003) was estimated in various studies to have noise shares of between 5 and 7% Wang et al. (2019); Reiss et al. (2020); Rücker and Akbik (2023). Other NER datasets have also been found to contain a share of incorrect labels, with OntoNotes4 estimated around 8% and WNUT-17 around 18% Wang et al. (2019); Huang et al. (2021).
@@
  • We comparatively evaluate current state-of-the-art approaches for noise-robust learning on NoiseBench, and experimentally establish upper bounds.
@@ 2 NoiseBench
Our benchmark is derived from a subset of the classic CoNLL-03 dataset for NER in English, annotated with entities belonging to four classes. We chose this dataset since it has been extensively studied in the field, allowing us to integrate various prior works. We derive a similar benchmark for NER in German, in Section 5 and Appendix A.
NoiseBench consists of the following parts: (1) A noise-free test split to evaluate trained models. (2) Seven variants of the training split, where six are annotated with different types of noise and one is without noise. Table 1 presents the quality of the six noisy variants w.r.t. the noise-free dataset.
@@
Table 1: Overview of the noisy training splits in NoiseBench. The table shows the noise level, the micro-averaged token-level F1 score (F1token), micro-averaged entity-level F1 (F1entity), the number of entities (Total), number of correct entities (Correct) and share of each error type: missing mentions (Missing (FN)), non-entity mentions (Non-entity (FP)), wrong type (Type) and partial matches (Partial). All metrics are in reference to the Clean split.
@@ 2.1.2 Expert Errors
Most machine learning datasets are created using manual annotation by domain experts that provide high-quality labels. However, errors have been found to occur even in expert annotation, affecting even well-known benchmarks, though usually with relatively low noise shares of under 10% Northcutt et al. (2021b); Song et al. (2022). To represent such noise, our benchmark includes a variant of the train split called Expert, which contains the original CoNLL-03 annotations. As Table 1 shows, this split has a noise share of 5.5% and is thus the split with lowest noise.
@@ 2.1.3 Crowdsourcing Errors
Crowdsourcing is a less costly alternative to expert annotation, but also more prone to annotation errors (Frenay and Verleysen, 2014). In order to create noisy variants of the train set representing real-world human errors, we utilize the crowdsourced labels by Rodrigues et al. (2014). This study involves 47 crowd workers labelling a subset of the English CoNLL-03 dataset, of around 400 news articles. They released their dataset and all annotations produced by each crowd worker. We selected only the sentences where the tokenization matched the Clean variant, resulting in 5,885 sentences.
@@ 2.1.4 Distant Supervision
One approach for labeling data without human participation is distant supervision Mintz et al. (2009), where entity mentions in target datasets are matched to entity types in knowledge bases (KBs).
We include a Distant noisy training variant in our benchmark, adapted from the annotations by Liang et al. (2020) [^2] Available under Apache 2.0 license that use the Wikidata corpus and gazetteers collected from multiple online sources as external knowledge bases. After initial POS tagging, the unlabeled sentences were matched with the knowledge bases. This process results in incomplete annotations due to limited coverage over entity types of KBs. This explains the rather high number of missing entities and the overall noise level (31.3%) of the Distant training variant, as shown in Table 1.
@@ 2.1.5 Weak Supervision
Another approach aimed at reducing manual annotation efforts is weak supervision. Here, labels are obtained using a number of “weak” supervision sources, such as heuristics or expression-based rules. Each weak source is typically specialized to detect only a subset of the correct labels.
We use the labels from the approach by Lison et al. (2020)2 to create our Weak label set. This covers 16 weak labeling sources Zhang et al. (2021b), including heuristics, gazetteers and predictions of NER models trained on other corpora. An example heuristic is detecting PER (person) entities using a pre-defined list of first names.
@@ 2.1.6 LLM Teacher Models
Our benchmark includes a noisy variant of the train split annotated by an LLM. This follows recent efforts that use LLMs for dataset generation Wang et al. (2023). Here, the main idea is to pass a description of the annotation task and target classes to an LLM, and provide sentences to label. LLMs are able to generate high quality labels for some tasks (e.g. sentiment classification) while for others (e.g. NER and question type categorization) the resulting labels are very noisy Golde et al. (2023).
@@ 2.2 Statistics
The table also shows the share of different error types. The errors are categorized into 4 main categories: missing mentions, non-entity mentions (false positives), incorrect entity type (where the boundary is correct, but type incorrect) and partial matches. Partial matches are special cases where the type is correct, but the mention boundary is only partially correct. Refer to Figure 1 for examples.
@@ 3 Comparing Real and Simulated Noise
3.1 Noise Simulation Methods
3.2 Experimental Setup
Validation splits. We evaluate the setting in which all available data to train a model is noisy, including the validation set. To obtain noisy validation sets for each of our 7 dataset variants, we split the noisy datasets into training and validation sets. All sentences from 66 news documents from 1996-08-24 comprise the validation set, which is around 17% of all sentences, and are left out from model training and used for hyperparameter tuning.
Baseline. For NER, as a baseline approach, we fine-tune an xlm-roberta-large transformer using the FLERT approach Schweter and Akbik (2021). It improves upon the regular fine-tuning setup by considering document-level features of a sentence to be tagged. We use a learning rate of 5e-6 and a batch size of 32, for a fixed number of 10 epochs, without early stopping. These parameters were obtained according to the performance on a noisy validation set, keeping in mind that larger batch sizes are more robust to noise Rolnick et al. (2017). We use entity-level micro F1-score.
3.3 Experiment 1: Impact of Label Noise on Test Performance
3.3.1 Results and Discussion
The results for uniform noise are shown in Appendix D. We initially established that uniform noise is less challenging for the model, so in the results for Experiments 1 and 2 we chose to focus solely on oracle class-dependent noise.
The main results from Experiment 1 for oracle class-dependent noise are shown in Table 2. Following are our main observations.
Label noise degrades performance. When we compare the test F1 scores of the real noisy variants with the average score of 93.99 achieved when training on the Clean variant, we see that model performance is affected by each noise type. As the noise levels increase, this impact becomes more pronounced, showing that the baseline model lacks robustness to real noise. Comparing the test F1 scores of the simulated noisy variants, we can see that noise of 5.9% in the training set results in a score comparable to training on the Clean variant. However, as simulated noise levels increase, the noise does degrade test set scores.
@@ 3.3.1 Results and Discussion
(a) Real noise
(b) Simulated class-dependent noise
Figure 2: F1 scores on different subsets of entities in the test set: all, seen (clean), seen (noisy) and unseen.
Models generalize to unseen entities well. Figure 2 shows F1 scores for seen and unseen entities separately, further distinguishing seen entities by whether their label in the training set was clean or noise-corrupted. In Figure 2(a) we see that for the Expert and Crowd++ noise types, the score on the seen (clean) and the unseen entities is comparable, which indicates the model has the ability to generalize to unseen entities well. As for the remaining training splits with noise levels of over 30%, noise also affects the performance on unseen entities.
Simulated noisy patterns are disregarded. For all real noise types, the score on the seen (noisy) entities is low. With simulated noise however, in Figure 2(b) we see that for Expert and Crowd++, the score on the seen-noisy entities and seen-clean entities is close. This means that at low noise levels, the models are able to disregard simulated noisy patterns and predict the same entities correctly when they appear in the test set.
@@
Figure 3: Comparison of model performance during extended training. The top row shows models fine-tuned on label sets with real noise, while the bottom row models fine-tuned on corresponding simulated (class-dependent) noisy labels. The plots are averages of 3 runs. The graphs for Crowd, Weak and LLM are shown in Appendix E.
@@ 3.4.1 Results and Discussion
In Figure 3 we show training curves from training with real and simulated variants of NoiseBench  for 3 noise types: Expert, Crowd++ and Distant. We plot two scores: the F1-score on the respective noisy variant of the training set, and the F1 score on the Clean variant of the training set. In all training curves, we can observe the memorization effect, with each model perfectly fitting the noisy data by the end of training and reaching an F1 score close to 1.
Delayed memorization of simulated noise. However, we note that with simulated noise (see Figure 3, 3, 3) this happens much later in the training process than with real noise. In addition, the training curves of simulated noise show a stage during the early epochs where the score on the clean labels is consistently higher than the score on the noisy labels. This confirms previous findings that the model is able to learn general patterns first, before starting to memorize the noise.
Immediate memorization of real noise. With real noise, this does not happen and the model starts fitting the noisy labels from the beginning (see Figure 3, 3, 3). As a result, the score on the clean labels is consistently lower than the score on the noisy labels, throughout the training run [^3] We confirm this finding for German in Appendix A.2.2..
Our experiments find that real noise does not display distinct generalization/memorization phases during training, and rather immediately begins with memorization [^4] We confirm this finding for a smaller model, as well as a randomly initialized model in Appendix E.. This makes intuitive sense, as real noise has underlying patterns that may be extracted during learning. This lends further evidence to the increased challenges and the need to evaluate noise-robust learning with real noise.
@@ 4.1 Compared Approaches
We surveyed current state-of-the-art methods for noise-robust NER and found that many approaches rely on the same underlying ideas for handling label noise. In the following, we group approaches by the underlying idea, select a state-of-the-art representative for each group and, if possible, derive an upper bound method for each group. For more details about the implementation of compared approaches refer to Appendix B.
@@ 4.1.1 Learning from a Clean Subset
The first family of approaches relies on utilizing the subset of each noisy dataset in which all labels are correct. One type of these approaches filters out all likely incorrect annotations and learns only from a clean subset. Another type derives confidence weights for each sample so that annotations judged to be of higher quality feature more during training.
As a representative of the former type of approaches targeting clean subsets of noisy datasets, we chose Confident Learning Northcutt et al. (2021a), while the latter type is represented by CrossWeigh Wang et al. (2019) and Learn-To-Reweight (L2R) Ren et al. (2018).
Upper bound: Oracle subset. To obtain an upper bound for this family of approaches, we use an oracle to select the subset of clean sentences from each of the noisy training splits in NoiseBench. We then use the baseline fine-tuning approach only on this subset, illustrating a best-case scenario.
@@ 4.1.2 Delaying Memorization
Another family of noise-robust learning approaches seeks to leverage the two phases of learning (generalization and memorization) we discussed in Section 3.4. They seek to either draw our the generalization phase or cease training before memorization begins. While our experiments indicate that these two phases do not exist for real noise, we nevertheless include this family of approaches in our evaluation since they are widely used. As representative of this class of approaches, we chose co-regularization (Zhou and Chen, 2021).
Upper bound: Oracle stopping. To obtain an upper bound for this family of approaches, we use a simple stopping criterion based on the score on the clean test set at the end of each epoch. We use the epoch of best generalization to report the final score. This simulates an ideal stopping.
@@ 4.1.3 Combined Approaches
While the approaches discussed so far each build on the individual ideas of identifying a clean subset or delaying memorization, many current approaches in fact combine multiple of such ideas in multi-stage pipelines Liang et al. (2020); Yu et al. (2021); Wang et al. (2022). As representative of such approaches, we evaluate BOND Liang et al. (2020) and meta self-refinement (MSR) Zhu et al. (2023a), both of which combine pseudo-labeling in a student-teacher setup and confidence-based sample selection.
]moved some implementation details to appendix
No upper bound for pseudo-labeling. We cannot derive a separate upper bound for pseudo-labeling, as the best case scenario here would mean that all noisy labels are replaced by correct labels, which is the same as training on the Clean dataset.
4.1.4 Additional Clean Data
We include a further upper bound for the scenario in which a small amount of high quality noise-free data is available. This is inspired by the extensive analysis of the use of clean validation data in Zhu et al. (2023b). Here, after first training on the noisy training set, they use a small clean dataset to continue fine-tuning the model. We include this upper bound to measure the accuracy gains that may be achieved if one were to invest effort in manually annotating additional noise-free data.